# X-ray Insights Unleashed: Pioneering the Enhancement of Multi-Label Long-Tail Data

### v1

@@ 1 Introduction
Chest X-ray (CXR), as the most common technique in the realm of medical imaging, plays a pivotal role in the diagnosis of thoracic diseases. Despite the rapidity and convenience of CXR imaging, the onus falls upon radiologists to manually scrutinize thousands of radiological samples daily, a task that is both arduous and burdensome. With the advent of deep learning paradigms and the burgeoning availability of CXR datasets [bustos2020padchest, johnson2019mimi, nguyen2022vindr, irvin2019chexpert, wang2017chestx], significant strides have been made in the realm of CXR diagnostics, particularly in disease detection [haque2024robust], automated report generation [liu2021auto], and CXR image synthesis [bluethgen2024vision].
@@
Figure 2: (a) The data distribution in each public dataset. (b) Diagram of lung disease entanglement. Each point in the figure is the center of the lesion annotation in the VinDr-CXR [nguyen2022vindr] dataset.
Researchers have endeavored to address the long-tail problem in medical imaging through non-generative methods, which require additional statistical support, such as the computation of data distribution statistics, or involve labor-intensive procedures, including fine-tuning and the manual delineation of head-tail class divisions. Typical strategies encompass data re-sampling and loss re-weighting. Data re-sampling methods [gupta2019lvis] demand pre-computed data distributions, potentially overfitting in tail classes and underfitting in head classes. Similarly, loss re-weighting approaches [cui2019class, kang2019decoupling, li2020overcoming] depend on these statistics to adjust weights, elevating the importance of tail classes while diminishing that of head classes at the class level. Both methods are heavily relied on complex computations of prior data, which may subsequently constrain the enhancement of performance for tail classes.
Recently, researchers have explored the potential of training a CXR image generator to facilitate data augmentation. As shown in Figure 1(a), the CXR generator produces images based on either textual report descriptions [bluethgen2024vision] or stochastic noise inputs [shentu2024cxr]. However, this approach encounters challenges in addressing tail classes, as the training data for the generative model also exhibits a long-tail distribution. It becomes evident that training a generator capable of reliably producing high-quality data for tail classes is an impractical expectation. Contrary to tail class data, the quantity of normal samples in clinical practice significantly outnumbers that of abnormal samples. This observation inspires us to consider:
@@
Intuitively, the proposition is to train a normal CXR generator that inpaints regions of easily detected lesions (e.g., head classes) to restore normal lung texture, while the remaining challenging-to-detect categories (e.g., tail classes) constitute a new training image. As depicted in Figure 1(b), we initially train a normal X-ray diffusion model and a lung lesion classifier. The classifier identifies the disease areas of head classes using a class activation map (CAM) [selvaraju2017grad]. The CAM map is then transformed into a mask and inputted into the pre-trained CXR diffusion model to generate a new image. Through this process, the head classes in the image are eliminated, while the tail classes persist and become new training data, thereby augmenting the number of tail samples. However, this process encounters several challenges: Disease Entanglement: The imaging characteristics of chest X-rays may cause different disease categories to become entangled (as shown in Figure 2(b)). When head and tail classes appear in the same area, inpainting the head class may lead to incorrect data generation. Domain Gap: Given that the CXR generator is trained on disparate datasets, the domain of the inpainted image may diverge from that of the original dataset, leading to a decline in model performance.
@@ 2.1 Lung Disease Recognition
With the development of deep learning, researchers have proposed many methods for automatic disease recognition in chest X-ray (CXR) images. Cervantes et al. [cervantes2021lime] proposed a simple and effective approach for COVID-19 screening by analyzing CXR images using CNNs. The study utilized four pre-trained CNN networks from ImageNet: ResNet50, VGG16, DenseNet201 and EfficientNetB3, for finetuning on a dataset containing CXR images of COVID-19 cases, healthy individuals, and patients with viral pneumonia. Priya et al. [mohanapriya2019lung] devised a lung diseases recognization system relying on deep convolutional neural networks, emphasizing the potential of CNNs in detecting lung abnormalities such as cancer. This study highlighted the importance of transfer learning, which enables leveraging pretrained CNN architectures for smaller datasets, thus contributing to the success of such studies. Abdul et al. [abdul2020automatic] introduced an automatic lung cancer detection and classification system designed to solve the challenge of early lung cancer detection using computed tomography images.
Afterward, some works applied Vision Transformer (ViT) [dosovitskiy2020image] to improve the visual feature extraction in the chest X-rays. Chen et al. [chen2024vision] analyzed and compared the performance of four diverse architectures: ViT, EfficientNet [tan2019efficientnet], MViT [li2022mvitv2], and EfficientViT [liu2023efficientvit], to explore the influence of varying the number of encoder blocks within the transformer models. Rocha et al. [rocha2024stern] proposed STERN, composed of an attention-driven spatial transformer module and a classifier, which is able to select the thoracic region of interest, transform the input image accordingly, and then classify it as normal or abnormal. Zhang et al. [zhang2024three] proposed a three-stage framework with knowledge transfer from adult chest X-rays to aid the diagnosis and interpretation of pediatric thorax diseases.
@@ 2.2 Long Tail Data Augmentation
Researchers have proposed many solutions for long-tail datasets, which can be divided into generative and non-generative. Representatives of non-generative solutions are data augmentation and loss reweighting. Galdran et al. [galdran2021balanced] proposed Balanced-MixUp, an extension of the MixUp [zhang2017mixup] regularization technique with class-balanced sampling, a general augmentation approach in the long-tail learning. Ju et al. [ju2021relational] classified tail classes into subsets based on prior information (clinical presentation and location) and used knowledge distillation to train a teacher model to enforce the student model to learn these prior knowledge. Zhang et al. [zhang2021mbnm] combined a feature memory module, resampling of tail classes, and a re-weighted loss function to improve the long-tail classification performance. Kang et al. [kang2019decoupling] proposed to decouple the learning procedure into representation learning and classification, and systematically explore how different balancing strategies affect for long-tailed recognition. Li et al. [li2020overcoming] proposed a novel balanced group softmax module for balancing the classifiers within the detection frameworks through group-wise training. Kozerawski et al. [kozerawski2020blt] proposed a data augmentation technique that compensates the imbalance of long-tailed classes by generating hard examples via gradient ascent techniques from existing tail-class training examples.
The main idea of the generative method is to increase the amount of data in the tail classes. Shen et al. [shen2023image] proposed a novel lung nodule synthesis framework based on image inpainting, which hierarchically decomposes the nodule synthesis procedure into shape generation, size modulation, and texture synthesis. Tang et al. [tang2021disentangled] proposed a deep disentangled generative model simultaneously generating abnormal disease residue maps and “radiorealistic” normal CXR images from an input abnormal CXR image. Luo et al. [luo2024measurement] devised a new diffusion synthesis framework, which introduces uncertainty guidance in each sampling step and design an uncertainty-guided diffusion models. Tang et al. [tang2024text] introduced a novel text-guided tail-class generation network, which first learns the feature representations of from each other via knowledge mutual distillation network, thereby enhancing the feature extraction capability of the tail-class model.
@@ 3.1.2 Lung Lesion Classifier Training on Abnormal X-rays
With the pre-trained normal X-ray generator established, the subsequent step requires localizing the head-class lesions to guide the inpainting process. To achieve this, we first train a lung lesion classifier on the collected abnormal dataset. We then employ Gradient-weighted Class Activation Mapping (Grad-CAM [selvaraju2017grad]) to extract the precise regions of interest corresponding to the head classes from this classifier. As illustrated in Figure 3, this pre-trained classifier and the associated Grad-CAM module constitute the second step of our pipeline, which is dedicated to lesion area extraction.
@@ 4.1 Datasets
Normal X-ray Dataset. We extract normal X-rays from the public datasets [nguyen2022vindr, irvin2019chexpert, wang2017chestx, johnson2019mimic, bustos2020padchest], to form a rich, multi-source normal X-ray dataset. Specifically, 215,652 normal X-ray images were collected from 7 publicly available datasets. In addition, we also collect a large number of normal X-rays (105,234) from our partner hospitals. The dataset details are given in Table 1.
@@
VinDr-CXR [nguyen2022vindr] 10606
CheXpert [irvin2019chexpert] 17002
NIH-CXR [wang2017chestx] 60362
MIMIC-CXR [johnson2019mimic] 81117
PadChest [bustos2020padchest] 36129
@@
Evaluation Dataset. We use two public chest X-ray datasets (MIMIC-CXR[johnson2019mimic] and CheXpert [irvin2019chexpert]) to evaluate the performance of our proposed method. Both datasets have the same 13 lesion categories and consistent long-tail distribution, as shown in Figure 2(a). The 13 lesion categories are Enlarged Cardiomediastinum (EC), Cardiomegaly (CA), Lung Opacity (LO), Lung Lesion (LL), Edema (ED), Consolidation (CO), Pneumonia (PA), Atelectasis (AT), Pneumothorax (PX), Pleural Effusion (PE), Pleural Other (PO), Fracture (FE), and Support Devices (SS). TThere are 150,904 samples in MIMIC-CXR and 166,458 samples in CheXpert. We randomly selected 80% of the dataset for training sets and the remaining 20% for testing.
@@ 4.3 Performance Analysis
Improvement in Different Classification Networks. To evaluate the effectiveness of our proposed normal data inpainting method, we apply this data augmentation on different classification networks (i.e., CNN-based networks, including ResNet [he2016deep], EfficientNet [tan2019efficientnet], Convnext [liu2022convnet] and Transformer-based networks, Swin Transformer [liu2021swin] and ViT [dosovitskiy2020image]), results are given in Table 2. From the table we can see that the performance of CNN-based networks is better than that of transformer-based networks, where EfficientNet shows the best performance on both datasets, 40.88% and 40.51%, respectively, while ViT has the worst performance, 34.01% and 31.97%, respectively. In terms of the performance improvement of the tail class, after the inpainting tail classes data was added to the network training, the performance of the tail classes was greatly improved. Although the performance of the head classes has variations, it is clinically acceptable to lose some of the performance of the head classes in exchange for a significant improvement in the performance of the tail classes. These results proved the effectiveness of our proposed data augmentation method of tail classes.
@@
Comparison to Caption-based Generation. To further demonstrate the effectiveness of our proposed normal X-ray inpainting method, we designed a variant of DiT (see Figure 5) by changing the original label input to the lesion category name. We used CLIP’s text encoder [radford2021learning] to encode the lesion name as the network input and generated tail-class images by inputting the name of the tail-class lesion. The experimental results are shown in Table 3. As seen from the table, although the caption-based method also improves the performance of tail classes, its improvement is significantly lower than that of our inpainting-based method. Additionally, in terms of head-class lesions performance, the caption-based method does not outperform our method. We speculate that this result is due to the long-tail problem in the training dataset, which affects the training of DiT and, consequently, the quality of the generated data. The experimental results show that when the data used to train the diffusion model exhibits a long-tail distribution, the quality of the generated images decreases, thereby affecting the performance improvement of the network. In contrast, our proposed diffusion model, trained on normal X-ray data, demonstrates significant performance improvements due to the large volume of available training data.
Ablation Study. In Table 4, we conducted ablation experiments to verify the effectiveness of the proposed progressive incremental learning strategy. Moreover, we introduced the focal loss [lin2017focal] to compare the performance improvement in tail classes. ResNet50 is taken as the verification network. From the table, we can see that although the classification accuracy achieved by the focal loss is higher than that of the cross-entropy loss for some head categories, it is lower than that of the cross-entropy loss in the tail categories. When PIL is not introduced into the network, the head classes will suffer from catastrophic forgetting due to the large amount of tail class data added to the training, which greatly reduces the performance of the head class. When PIL is added to the network training, the performance of the head classes is retained, and even improved for a few head classes. These results validated the effectiveness of our proposed PIL strategy.

### v2

@@ 1 Introduction
Chest X-ray (CXR), as the most common technique in the realm of medical imaging, plays a pivotal role in the diagnosis of thoracic diseases. Despite the rapidity and convenience of CXR imaging, the onus falls upon radiologists to manually scrutinize thousands of radiological samples daily, a task that is both arduous and burdensome. With the advent of deep learning paradigms and the burgeoning availability of CXR datasets [3, 14, 27, 12, 36], significant strides have been made in the realm of CXR diagnostics, particularly in disease detection [10], automated report generation [21], and CXR image synthesis [2].
@@
Figure 2: (a) The data distribution in each public dataset. (b) Diagram of lung disease entanglement. Each point in the figure is the center of the lesion annotation in the VinDr-CXR [27] dataset.
Researchers have endeavored to address the long-tail problem in medical imaging through non-generative methods, which require additional statistical support, such as the computation of data distribution statistics, or involve labor-intensive procedures, including fine-tuning and the manual delineation of head-tail class divisions. Typical strategies encompass data re-sampling and loss re-weighting. Data re-sampling methods [9] demand pre-computed data distributions, potentially overfitting in tail classes and underfitting in head classes. Similarly, loss re-weighting approaches [6, 16, 19] depend on these statistics to adjust weights, elevating the importance of tail classes while diminishing that of head classes at the class level. Both methods are heavily relied on complex computations of prior data, which may subsequently constrain the enhancement of performance for tail classes.
Recently, researchers have explored the potential of training a CXR image generator to facilitate data augmentation. As shown in Figure 1(a), the CXR generator produces images based on either textual report descriptions [2] or stochastic noise inputs [32]. However, this approach encounters challenges in addressing tail classes, as the training data for the generative model also exhibits a long-tail distribution. It becomes evident that training a generator capable of reliably producing high-quality data for tail classes is an impractical expectation. Contrary to tail class data, the quantity of normal samples in clinical practice significantly outnumbers that of abnormal samples. This observation inspires us to consider:
@@
Intuitively, the proposition is to train a normal CXR generator that inpaints regions of easily detected lesions (e.g., head classes) to restore normal lung texture, while the remaining challenging-to-detect categories (e.g., tail classes) constitute a new training image. As depicted in Figure 1(b), we initially train a normal X-ray diffusion model and a lung lesion classifier. The classifier identifies the disease areas of head classes using a class activation map (CAM) [30]. The CAM map is then transformed into a mask and inputted into the pre-trained CXR diffusion model to generate a new image. Through this process, the head classes in the image are eliminated, while the tail classes persist and become new training data, thereby augmenting the number of tail samples. However, this process encounters several challenges: Disease Entanglement: The imaging characteristics of chest X-rays may cause different disease categories to become entangled (as shown in Figure 2(b)). When head and tail classes appear in the same area, inpainting the head class may lead to incorrect data generation. Domain Gap: Given that the CXR generator is trained on disparate datasets, the domain of the inpainted image may diverge from that of the original dataset, leading to a decline in model performance.
@@ 2.1 Lung Disease Recognition
With the development of deep learning, researchers have proposed many methods for automatic disease recognition in chest X-ray (CXR) images. Cervantes et al. [4] proposed a simple and effective approach for COVID-19 screening by analyzing CXR images using CNNs. The study utilized four pre-trained CNN networks from ImageNet: ResNet50, VGG16, DenseNet201 and EfficientNetB3, for finetuning on a dataset containing CXR images of COVID-19 cases, healthy individuals, and patients with viral pneumonia. Priya et al. [26] devised a lung diseases recognization system relying on deep convolutional neural networks, emphasizing the potential of CNNs in detecting lung abnormalities such as cancer. This study highlighted the importance of transfer learning, which enables leveraging pretrained CNN architectures for smaller datasets, thus contributing to the success of such studies. Abdul et al. [1] introduced an automatic lung cancer detection and classification system designed to solve the challenge of early lung cancer detection using computed tomography images.
Afterward, some works applied Vision Transformer (ViT) [7] to improve the visual feature extraction in the chest X-rays. Chen et al. [5] analyzed and compared the performance of four diverse architectures: ViT, EfficientNet [33], MViT [18], and EfficientViT [22], to explore the influence of varying the number of encoder blocks within the transformer models. Rocha et al. [29] proposed STERN, composed of an attention-driven spatial transformer module and a classifier, which is able to select the thoracic region of interest, transform the input image accordingly, and then classify it as normal or abnormal. Zhang et al. [39] proposed a three-stage framework with knowledge transfer from adult chest X-rays to aid the diagnosis and interpretation of pediatric thorax diseases.
@@ 2.2 Long Tail Data Augmentation
Researchers have proposed many solutions for long-tail datasets, which can be divided into generative and non-generative. Representatives of non-generative solutions are data augmentation and loss reweighting. Galdran et al. [8] proposed Balanced-MixUp, an extension of the MixUp [37] regularization technique with class-balanced sampling, a general augmentation approach in the long-tail learning. Ju et al. [15] classified tail classes into subsets based on prior information (clinical presentation and location) and used knowledge distillation to train a teacher model to enforce the student model to learn these prior knowledge. Zhang et al. [38] combined a feature memory module, resampling of tail classes, and a re-weighted loss function to improve the long-tail classification performance. Kang et al. [16] proposed to decouple the learning procedure into representation learning and classification, and systematically explore how different balancing strategies affect for long-tailed recognition. Li et al. [19] proposed a novel balanced group softmax module for balancing the classifiers within the detection frameworks through group-wise training. Kozerawski et al. [17] proposed a data augmentation technique that compensates the imbalance of long-tailed classes by generating hard examples via gradient ascent techniques from existing tail-class training examples.
The main idea of the generative method is to increase the amount of data in the tail classes. Shen et al. [31] proposed a novel lung nodule synthesis framework based on image inpainting, which hierarchically decomposes the nodule synthesis procedure into shape generation, size modulation, and texture synthesis. Tang et al. [35] proposed a deep disentangled generative model simultaneously generating abnormal disease residue maps and “radiorealistic” normal CXR images from an input abnormal CXR image. Luo et al. [25] devised a new diffusion synthesis framework, which introduces uncertainty guidance in each sampling step and design an uncertainty-guided diffusion models. Tang et al. [34] introduced a novel text-guided tail-class generation network, which first learns the feature representations of from each other via knowledge mutual distillation network, thereby enhancing the feature extraction capability of the tail-class model.
@@ 3.1.2 Lung Lesion Classifier Training on Abnormal X-rays
With the pre-trained normal X-ray generator established, the subsequent step requires localizing the head-class lesions to guide the inpainting process. To achieve this, we first train a lung lesion classifier on the collected abnormal dataset. We then employ Gradient-weighted Class Activation Mapping (Grad-CAM [30]) to extract the precise regions of interest corresponding to the head classes from this classifier. As illustrated in Figure 3, this pre-trained classifier and the associated Grad-CAM module constitute the second step of our pipeline, which is dedicated to lesion area extraction.
@@ 4.1 Datasets
Normal X-ray Dataset. We extract normal X-rays from the public datasets [27, 12, 36, 13, 3], to form a rich, multi-source normal X-ray dataset. Specifically, 215,652 normal X-ray images were collected from 7 publicly available datasets. In addition, we also collect a large number of normal X-rays (105,234) from our partner hospitals. The dataset details are given in Table 1.
@@
VinDr-CXR [27] 10606
CheXpert [12] 17002
NIH-CXR [36] 60362
MIMIC-CXR [13] 81117
PadChest [3] 36129
@@
Evaluation Dataset. We use two public chest X-ray datasets (MIMIC-CXR[13] and CheXpert [12]) to evaluate the performance of our proposed method. Both datasets have the same 13 lesion categories and consistent long-tail distribution, as shown in Figure 2(a). The 13 lesion categories are Enlarged Cardiomediastinum (EC), Cardiomegaly (CA), Lung Opacity (LO), Lung Lesion (LL), Edema (ED), Consolidation (CO), Pneumonia (PA), Atelectasis (AT), Pneumothorax (PX), Pleural Effusion (PE), Pleural Other (PO), Fracture (FE), and Support Devices (SS). TThere are 150,904 samples in MIMIC-CXR and 166,458 samples in CheXpert. We randomly selected 80% of the dataset for training sets and the remaining 20% for testing.
@@ 4.3 Performance Analysis
Improvement in Different Classification Networks. To evaluate the effectiveness of our proposed normal data inpainting method, we apply this data augmentation on different classification networks (i.e., CNN-based networks, including ResNet [11], EfficientNet [33], Convnext [24] and Transformer-based networks, Swin Transformer [23] and ViT [7]), results are given in Table 2. From the table we can see that the performance of CNN-based networks is better than that of transformer-based networks, where EfficientNet shows the best performance on both datasets, 40.88% and 40.51%, respectively, while ViT has the worst performance, 34.01% and 31.97%, respectively. In terms of the performance improvement of the tail class, after the inpainting tail classes data was added to the network training, the performance of the tail classes was greatly improved. Although the performance of the head classes has variations, it is clinically acceptable to lose some of the performance of the head classes in exchange for a significant improvement in the performance of the tail classes. These results proved the effectiveness of our proposed data augmentation method of tail classes.
@@
Comparison to Caption-based Generation. To further demonstrate the effectiveness of our proposed normal X-ray inpainting method, we designed a variant of DiT (see Figure 5) by changing the original label input to the lesion category name. We used CLIP’s text encoder [28] to encode the lesion name as the network input and generated tail-class images by inputting the name of the tail-class lesion. The experimental results are shown in Table 3. As seen from the table, although the caption-based method also improves the performance of tail classes, its improvement is significantly lower than that of our inpainting-based method. Additionally, in terms of head-class lesions performance, the caption-based method does not outperform our method. We speculate that this result is due to the long-tail problem in the training dataset, which affects the training of DiT and, consequently, the quality of the generated data. The experimental results show that when the data used to train the diffusion model exhibits a long-tail distribution, the quality of the generated images decreases, thereby affecting the performance improvement of the network. In contrast, our proposed diffusion model, trained on normal X-ray data, demonstrates significant performance improvements due to the large volume of available training data.
Ablation Study. In Table 4, we conducted ablation experiments to verify the effectiveness of the proposed progressive incremental learning strategy. Moreover, we introduced the focal loss [20] to compare the performance improvement in tail classes. ResNet50 is taken as the verification network. From the table, we can see that although the classification accuracy achieved by the focal loss is higher than that of the cross-entropy loss for some head categories, it is lower than that of the cross-entropy loss in the tail categories. When PIL is not introduced into the network, the head classes will suffer from catastrophic forgetting due to the large amount of tail class data added to the training, which greatly reduces the performance of the head class. When PIL is added to the network training, the performance of the head classes is retained, and even improved for a few head classes. These results validated the effectiveness of our proposed PIL strategy.